OpenStudy (anonymous):

Find four consecutive even integers such that twice the sum of the second and third exceeds 3 times the first by 34 Find four consecutive even integers such that twice the sum of the second and third exceeds 3 times the first by 34 @Mathematics

OpenStudy (anonymous):

22, 24, 26, 28

OpenStudy (anonymous):

Let x be the the third of 4 consecutive even number respectively. Solve the following for x: 2 ((x - 2) + x) = 3 (x - 4) + 34 -4 + 4 x = 22 + 3 x x = 26, {22, 24, 26, 28}